Consider four different stocks, all of which have a required return of 18.25 percent and a most recent dividend of $3.10 per share. Stocks W, X, and Y are expected to maintain constant growth rates in dividends for the foreseeable future of 11 percent, 0 percent, and –5.5 percent per year, respectively. Stock Z is a growth stock that will increase its dividend by 20.25 percent for the next two years and then maintain a constant 13 percent growth rate, thereafter.

a. What is the dividend yield for each of these four stocks? (Do not round intermediate calculations and enter your answers as a percent rounded to 2 decimal places, e.g., 32.16.)
b. What is the expected capital gains yield for each of these four stocks? (A negative answer should be indicated by a minus sign. Leave no cells blank - be certain to enter "0" wherever required. Do not round intermediate calculations and enter your answers as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1

A C D E LL G 1 B a) Stock W 2 3 4 5 D1 3.44 3.10 2.93 Price 47.46 16.99 12.33 Х Dividend yield 7.25% 18.25% 23.75% 4.95% Y LC

EXCEL FORMULA:

A B C D E a) 1 2 Dividend yield 3 Stock W X 4 D1 =3.1*(1+11%) |=3.1*(1+0%) =3.1*(1-5.5%) =3.1*(1+20.25%) Price =C3/(18.25%-11